A person sitting on the tall building is dropping balls at regular intervals of one second. Find the position of the 3,4,5th ball when the 6th ball is being dropped

Grade:11

1 Answers

Adarsh
6 years ago
Dear Swati,
The position of 3rd 4th and 5th ball will be equal to the distance covered by ball in 1 , 2 and  3 second.
So distance covered in 1st second = 0.t+1/2.a.t2=1/2.10.1=5m
Distance covered in 2 second =0.t+1/2.a.t2=1/2.10.2.2=20m
Distance covered in 3 second =0.t+1/2.a.t2=1/2.10.3.3=45m
Hence distance covered by 3rd,4th and 5th ball is 5m , 20m and 45m.
 
REGARDS
Adarsh